# Distinguishing the origin of eccentric black-hole mergers with gravitational-wave spin measurements

### v1

@@ 3 Results
Figure 2: Bayes factor distributions that compare the mutually aligned and uncorrelated model. The red histogram assumes GW detections of systems with mutually aligned spins and indicates the confidence for preferring the aligned model over the uncorrelated model. Conversely, the teal histogram assumes GW detections of systems with uncorrelated spins and indicates the confidence for preferring the uncorrelated spin model over the mutually aligned spin model. Thus, in either cases a large Bayes factor ($\log_{10}\mathcal{B}\gtrsim 1/2$) indicates that we would correctly infer the spin properties of the underlying population. Solid and dashed vertical lines indicate the medians and 5th/95th percentiles of the distributions, respectively. We assume an eccentric BBH fraction of $\xi_{\rm ecc}=0.05$, a measurement uncertainty for $\chi_{\rm eff}$ of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE, and vary between each panel the total number of detections between $2,100$ (LVK O5), $10^{4}$ (A#), and $10^{5}$ (ET/CE).
@@
Figure 2 shows the Bayes factor distributions for the three detector configurations LVK O5 (right panel), A# (middle panel), and ET/CE (right panel), for our default assumptions $\xi_{\rm ecc}=0.05$ and $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE. If the eccentric BBH mergers form with mutually aligned spins (“Mutually aligned data”, red histograms), the Bayes factors for statistically preferring the mutually aligned spin model over the uncorrelated spin model range from $\log_{10}\mathcal{B}=0.37^{+1.14}_{-0.99}$ (LVK O5), $\log_{10}\mathcal{B}=1.84^{+2.77}_{-2.55}$ (A#), to $\log_{10}\mathcal{B}=111^{+87}_{-84}$ (ET/CE), where uncertainties refer to the 5th/95th percentiles. Similarly, for eccentric BBH mergers with uncorrelated spins (“Uncorrelated data”, teal histograms), we find Bayes factors for recovering the uncorrelated spin model over the mutually aligned spin model ranging from $\log_{10}\mathcal{B}=0.36^{+0.83}_{-0.97}$ (LVK O5), $\log_{10}\mathcal{B}=1.61^{+2.05}_{-2.27}$ (A#), to $\log_{10}\mathcal{B}=95^{+66}_{-62}$ (ET/CE). We consider $\log_{10}\mathcal{B}\gtrsim 1/2$ (non-hatched region) as a minimal threshold for substantial evidence to prefer the correct model. This is the case for $\sim 40\,\%$ of the realizations in LVK O5, $\sim 80\,\%$ for A#, and $\sim 98\,\%$ for ET/CE. Meanwhile, obtaining substantial evidence for preferring the incorrect model ($\log_{10}\mathcal{B}\lesssim-1/2$) happens for $\sim 7\,\%$ of the realizations in LVK O5, $\sim 0.06\,\%$ for A#, and $\sim 0.01\,\%$ for ET/CE. Thus, in all detector configurations, the probability of preferring the incorrect model is low. However, only with A# and ET/CE do we expect a large probability to infer the correct model at high significance, and thus to be able to statistically distinguish the dynamical origin of the eccentric BBHs. For LVK O5, it is most likely that the data will show no strong preference for either model.
@@
In Fig. 5, we vary the eccentric merger fraction $\xi_{\rm ecc}$ (while keeping our default measurement uncertainty of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE). For LVK O5, A#, and ET/CE, we find a minimum eccentric merger fraction of $\xi_{\rm eff}\gtrsim 9\,\%$, $1\,\%$, and $0.03\,\%$, respectively, is needed so that the median of the Bayes factor distributions exceeds $\log_{10}\mathcal{B}\approx 1/2$. The 5th percentile would exceed $\log_{10}\mathcal{B}\approx 1/2$ if $\xi_{\rm eff}\gtrsim 21\,\%$ (A#) and $0.4\,\%$ (ET/CE), respectively, whereas the 5th percentile remains below $\log_{10}\mathcal{B}\approx 1/2$ for LVK O5.
In Fig. 5, we vary the measurement uncertainty $\sigma_{\chi_{\rm eff}}$ (while keeping our default eccentric merger fraction $\xi_{\rm ecc}=0.05$). For LVK O5, A#, and ET/CE, we find that it must not be greater than $\sigma_{\chi_{\rm eff}}\approx 0.16$, $0.29$, and $0.47$, respectively, for the median to exceed $\log_{10}\mathcal{B}\approx 1/2$. The 5th percentile would $\log_{10}\mathcal{B}\approx 1/2$ if $\sigma_{\chi_{\rm eff}}\lesssim 0.07$ (A#) and $0.25$ (ET/CE), respectively.
Figure 3: Percentiles of the Bayes factor distributions as a function of the eccentric BBH fraction $\xi_{\rm ecc}$. We assume that $\chi_{\rm eff}$ will be measured with an uncertainty of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE. As in Fig. 2, solid and dashed lines indicate the median and 5th/95th percentiles, respectively. The vertical yellow dotted lines indicate the eccentric fraction of events identified by Gupte et al. (2024) (3 of 57 analysed events; “Gupte+”) and by Romero-Shaw et al. (2022) (4 of 62 analysed events; “RS+”), respectively.
@@
Figure 5: Percentiles of the Bayes factor distributions that compare the uncorrelated model (teal) to a tilted model (red), where we allow the spins to be tilted with respect to one another by $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})$. We assume an eccentric BBH fraction of $\xi_{\rm ecc}=0.05$ and a measurement uncertainty for $\chi_{\rm eff}$ of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE.
@@
Above, we assumed that in the mutually aligned model the two spins are perfectly aligned to one another, i.e., $\boldsymbol{\hat{\chi}}_{1}=\boldsymbol{\hat{\chi}}_{2}$. Generally, mutually aligned spins can be the result of massive binary stars whose stellar spins synchronise due to tides or other stellar binary interactions and whose black hole remnants inherit the alignment before driven to an eccentric merger due to the environment (e.g., tertiary companions in hierarchical triples or AGNs). However, some misalignment might occur if, e.g., the stars are born with uncorrelated spins or misalign during a mass-transfer episode (Stegmann & Antonini, 2021b) but tides are too inefficient to align them, the black holes form with spins which do not inherit the orientation from the rotation axes of the progenitor stars (Baibhav & Kalogera, 2024), or supernova kicks tilt the direction of the orbital angular momentum (Kalogera, 2000; Gerosa et al., 2018; Steinle & Kesden, 2021). Thus, we investigate another “tilted model” where, instead of setting $\boldsymbol{\hat{\chi}}_{1}=\boldsymbol{\hat{\chi}}_{2}$, we sample $\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2}$ uniformly between one and a given minimum value $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\in[-1,1)$ and assume a random azimuthal angle of $\boldsymbol{\hat{\chi}}_{2}$ about $\boldsymbol{\hat{\chi}}_{1}$. In Fig. 5, we assume that the eccentric BBH mergers form with tilted spins (“Tilted data”) with some minimum value $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})$ shown on the x-axis (with our default assumptions $\xi_{\rm ecc}=0.05$ and $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE). Red and teal percentiles of $\log_{10}\mathcal{B}$ show how well we could recover the tilted model over the aligned and uncorrelated models, respectively. Note that for $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})=-1$ and $+1$ the tilted model is equivalent to the uncorrelated and mutually aligned model, respectively. Thus, the test against the mutually aligned model (red) converges to $\log_{10}\mathcal{B}=0$ as $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\rightarrow 1$, and vice-versa the test against the uncorrelated model (teal) goes to zero as $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\rightarrow-1$. Crucially, the tilted model will be statistically preferred against the uncorrelated model (teal) even if we allow large spin-spin misalignments, e.g., the teal median is $\log_{10}\mathcal{B}\gtrsim 1/2$ for $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\gtrsim 0$ for A# and for $\gtrsim-0.9$ for ET/CE, respectively. Astrophysically, that means that we would be able to distinguish formation channels in which the spin direction would fully randomise, e.g., clusters, from those that tend to align the spins, e.g., hierarchical triples, even if we allow significant tilt angles between them.
4 Conclusion
@@
Specifically, we investigated if $\chi_{\rm eff}$ measurements of eccentric BBH mergers can be used to distinguish between models with mutually aligned and randomly oriented spins, which would help to identify the dynamical formation channel through which they must have formed. For our default assumptions, corresponding to an eccentric merger fraction $\xi_{\rm ecc}=0.05$ and measurement uncertainty $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE, we find a $\sim 40\,\%$ chance to confidently distinguish both models after LVK O5, $\sim 80\,\%$ for LIGO A#, and $\sim 98\,\%$ for ET and CE (Fig. 2). These results depend significantly on $\xi_{\rm ecc}$ and the future values of $\sigma_{\rm eff}$, both of which are uncertain, and may correlate. Parameter estimation studies using eccentric waveform models have shown correlations between $\chi_{\mathrm{eff}}$ and eccentricity, which could cause wider uncertainties on $\chi_{\mathrm{eff}}$ if detected eccentricity is poorly constrained (O’Shea & Kumar, 2023; Divyajyoti et al., 2024b); however, we note that since future ground-based GW detectors will be able to constrain eccentricities about two orders of magnitude lower than current detectors (Saini, 2024), this effect is likely to be small.
@@
Finally, we note that our analysis is simplistic in two ways. First, we have not included any non-eccentric detections to distinguish between formation channels of BBH mergers. Since current models predict $\xi_{\rm eff}\lesssim\mathcal{O}(0.1)$, a given number of eccentric mergers necessarily implies there is a much larger number of circularised mergers originating from the same channel. Thus, any viable formation channel must not only recover the spin-orbit properties of eccentric mergers but also be consistent with the rest of the population, where multiple channels (including isolated binary star evolution) might contribute (Zevin et al., 2021b). In turn, focussing on the eccentric mergers has the advantage that we know they must originate from a dynamical or triple channel and that these channels make distinct predictions for the spin-orbit properties. Second, we did not consider the possibility that multiple channels contribute to the formation of eccentric mergers. We expect this to influence our analysis only if they contribute at about the same order of magnitude, in which case in step (ii) of Sec. 2 a mixture model must be constructed that contains a fraction $10<\beta<$ 1 of mutually aligned mergers and a fraction $1-\beta$ of uncorrelated ones.

### v2

@@ III Results
Figure 2: Bayes factor distributions that compare the mutually aligned and uncorrelated model. The maroon histogram assumes GW detections of systems with mutually aligned spins and indicates the confidence for preferring the aligned model over the uncorrelated model. Conversely, the teal histogram assumes GW detections of systems with uncorrelated spins and indicates the confidence for preferring the uncorrelated spin model over the mutually aligned spin model. Thus, in either cases a large Bayes factor ($\log_{10}\mathcal{B}\gtrsim 1.3$) indicates that we would correctly infer the spin properties of the underlying population. Solid and dashed vertical lines indicate the medians and 5th/95th percentiles of the distributions, respectively. We assume an eccentric BBH fraction of $\xi_{\rm ecc}=0.05$, a measurement uncertainty for $\chi_{\rm eff}$ of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE, and vary between each panel the total number of detections between $2,100$ (LVK O5), $10^{4}$ (A#), and $10^{5}$ (ET/CE).
@@
Figure 2 shows the Bayes factor distributions for the three detector configurations LVK O5 (right panel), A# (middle panel), and ET/CE (right panel), for our default assumptions $\xi_{\rm ecc}=0.05$ and $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE. If the eccentric BBH mergers form with mutually aligned spins (“Mutually aligned data”, maroon histograms), the Bayes factors for statistically preferring the mutually aligned spin model over the uncorrelated spin model range from $\log_{10}\mathcal{B}=0.37^{+1.14}_{-0.99}$ (LVK O5), $\log_{10}\mathcal{B}=1.84^{+2.77}_{-2.55}$ (A#), to $\log_{10}\mathcal{B}=111^{+87}_{-84}$ (ET/CE), where uncertainties refer to the 5th/95th percentiles. Similarly, for eccentric BBH mergers with uncorrelated spins (“Uncorrelated data”, teal histograms), we find Bayes factors for recovering the uncorrelated spin model over the mutually aligned spin model ranging from $\log_{10}\mathcal{B}=0.36^{+0.83}_{-0.97}$ (LVK O5), $\log_{10}\mathcal{B}=1.61^{+2.05}_{-2.27}$ (A#), to $\log_{10}\mathcal{B}=95^{+66}_{-62}$ (ET/CE). We consider $\log_{10}\mathcal{B}\gtrsim 1.3$ (non-hatched region) as a minimal threshold for substantial evidence to prefer the correct model. This is the case for $\sim 9\,\%$ of the realizations in LVK O5, $\sim 63\,\%$ for A#, and $\sim 98\,\%$ for ET/CE. Meanwhile, obtaining substantial evidence for preferring the incorrect model ($\log_{10}\mathcal{B}\lesssim-1.3$) happens for $\sim 1\,\%$ of the realizations in all detectors. Thus, in all detector configurations, the probability of preferring the incorrect model is low. However, only with A# and ET/CE do we expect a large probability to infer the correct model at high significance, and thus to be able to statistically distinguish the dynamical origin of the eccentric BBHs. For LVK O5, it is most likely that the data will show no strong preference for either model.
@@
In Fig. 5, we vary the eccentric merger fraction $\xi_{\rm ecc}$ (while keeping our default measurement uncertainty of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE). For LVK O5, A#, and ET/CE, we find a minimum eccentric merger fraction of $\xi_{\rm eff}\gtrsim 21\,\%$, $5\,\%$, and $0.08\,\%$, respectively, is needed so that the median of the Bayes factor distributions exceeds $\log_{10}\mathcal{B}\approx 1.3$. The 5th percentile would exceed $\log_{10}\mathcal{B}\approx 1.3$ if $\xi_{\rm eff}\gtrsim 37\,\%$ (A#) and $0.6\,\%$ (ET/CE), respectively, whereas the 5th percentile remains below $\log_{10}\mathcal{B}\approx 1.3$ for LVK O5.
In Fig. 5, we vary the measurement uncertainty $\sigma_{\chi_{\rm eff}}$ (while keeping our default eccentric merger fraction $\xi_{\rm ecc}=0.05$). For LVK O5, A#, and ET/CE, we find that it must not be greater than $\sigma_{\chi_{\rm eff}}\approx 0.09$, $0.21$, and $0.40$, respectively, for the median to exceed $\log_{10}\mathcal{B}\approx 1.3$. The 5th percentile would $\log_{10}\mathcal{B}\approx 1.3$ if $\sigma_{\chi_{\rm eff}}\lesssim 0.03$ (A#) and $0.21$ (ET/CE), respectively. Thus, we also repeated our entire analysis by conservatively assuming that the measurement uncertainty $\sigma_{\chi_{\rm eff}}$ of eccentric sources with ET/CE does not improve but will be the same as for LVK O5 and A#, i.e., $\sigma_{\chi_{\rm eff}}=0.2$ (instead of our default assumption of 0.02). We find that this only marginally affects our conclusions, demonstrating that it is primarily the large number of detections with ET/CE that will allow distinguishing between both models. In particular, we find that still $\sim 96\,\%$ of the realizations with $\xi_{\rm ecc}=0.05$ will have $\log_{10}\mathcal{B}\gtrsim 1.3$ (compared to the $\sim 98\,\%$ for $\sigma_{\chi_{\rm eff}}=0.02$ and $\sim 0.02\,\%$ have $\log_{10}\mathcal{B}\lesssim 1.3$ (instead of $\sim 0.01\,\%$).
Figure 3: Percentiles of the Bayes factor distributions as a function of the eccentric BBH fraction $\xi_{\rm ecc}$. We assume that $\chi_{\rm eff}$ will be measured with an uncertainty of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE. As in Fig. 2, solid and dashed lines indicate the median and 5th/95th percentiles, respectively. The vertical yellow dotted lines indicate the eccentric fraction of events identified by Gupte et al. (2024a) (3 of 57 analysed events; “Gupte+”) and by Romero-Shaw et al. (2022) (4 of 62 analysed events; “RS+”), respectively.
@@
Figure 5: Percentiles of the Bayes factor distributions that compare the uncorrelated model (teal) to a tilted model (maroon), where we allow the spins to be tilted with respect to one another by $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})$. We assume an eccentric BBH fraction of $\xi_{\rm ecc}=0.05$ and a measurement uncertainty for $\chi_{\rm eff}$ of $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE.
@@
Above, we assumed that in the mutually aligned model the two spins are perfectly aligned to one another, i.e., $\boldsymbol{\hat{\chi}}_{1}=\boldsymbol{\hat{\chi}}_{2}$. Generally, mutually aligned spins can be the result of massive binary stars whose stellar spins synchronise due to tides or other stellar binary interactions and whose black hole remnants inherit the alignment before driven to an eccentric merger due to the environment (e.g., tertiary companions in hierarchical triples or AGNs). However, some misalignment might occur if, e.g., the stars are born with uncorrelated spins or misalign during a mass-transfer episode (Stegmann and Antonini, 2021) but tides are too inefficient to align them, the black holes form with spins which do not inherit the orientation from the rotation axes of the progenitor stars (Baibhav and Kalogera, 2024a), or supernova kicks tilt the direction of the orbital angular momentum (Kalogera, 2000; Gerosa et al., 2018; Steinle and Kesden, 2021). Thus, we investigate another “tilted model” where, instead of setting $\boldsymbol{\hat{\chi}}_{1}=\boldsymbol{\hat{\chi}}_{2}$, we sample $\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2}$ uniformly between one and a given minimum value $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\in[-1,1)$ and assume a random azimuthal angle of $\boldsymbol{\hat{\chi}}_{2}$ about $\boldsymbol{\hat{\chi}}_{1}$. In Fig. 5, we assume that the eccentric BBH mergers form with tilted spins (“Tilted data”) with some minimum value $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})$ shown on the x-axis (with our default assumptions $\xi_{\rm ecc}=0.05$ and $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE). Maroon and teal percentiles of $\log_{10}\mathcal{B}$ show how well we could recover the tilted model over the aligned and uncorrelated models, respectively. Note that for $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})=-1$ and +1+1 the tilted model is equivalent to the uncorrelated and mutually aligned model, respectively. Thus, the test against the mutually aligned model (maroon) converges to $\log_{10}\mathcal{B}=0$ as $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\rightarrow 1$, and vice-versa the test against the uncorrelated model (teal) goes to zero as $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\rightarrow-1$. Crucially, the tilted model will be statistically preferred against the uncorrelated model (teal) even if we allow large spin-spin misalignments, e.g., the teal median is $\log_{10}\mathcal{B}\gtrsim 1.3$ for $\min(\boldsymbol{\hat{\chi}}_{1}\cdot\boldsymbol{\hat{\chi}}_{2})\gtrsim 0.7$ for A# and for $\gtrsim-0.9$ for ET/CE, respectively. Astrophysically, that means that we would be able to distinguish formation channels in which the spin direction would fully randomise, e.g., clusters, from those that tend to align the spins, e.g., hierarchical triples, even if we allow significant tilt angles between them.
IV Conclusion
@@
Specifically, we investigated if $\chi_{\rm eff}$ measurements of eccentric BBH mergers can be used to distinguish between models with mutually aligned and randomly oriented spins, which would help to identify the dynamical formation channel through which they must have formed. For our default assumptions, corresponding to an eccentric merger fraction $\xi_{\rm ecc}=0.05$ and measurement uncertainty $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# and $\sigma_{\chi_{\rm eff}}=0.02$ for ET/CE, we find an $\sim 9\,\%$ chance to confidently distinguish both models after LVK O5, $\sim 63\,\%$ for LIGO A#, and $\sim 98\,\%$ for ET and CE (Fig. 2). These results depend significantly on $\xi_{\rm ecc}$ and the future values of $\sigma_{\rm eff}$, both of which are uncertain, and may correlate. Parameter estimation studies using eccentric waveform models have shown correlations between $\chi_{\mathrm{eff}}$ and eccentricity, which could cause wider uncertainties on $\chi_{\mathrm{eff}}$ if detected eccentricity is poorly constrained (O’Shea and Kumar, 2023; Divyajyoti et al., 2024); however, we note that since future ground-based GW detectors will be able to constrain eccentricities about two orders of magnitude lower than current detectors (Saini, 2024), this effect is likely to be small.
@@
Finally, we note that our analysis is simplistic in four ways. First, we have not included any non-eccentric detections to distinguish between formation channels of BBH mergers. Since current models predict $\xi_{\rm eff}\lesssim\mathcal{O}(0.1)$, a given number of eccentric mergers necessarily implies there is a much larger number of circularised mergers originating from the same channel. Thus, any viable formation channel must not only recover the spin-orbit properties of eccentric mergers but also be consistent with the rest of the population, where multiple channels (including isolated binary star evolution) might contribute (Zevin et al., 2021a). In turn, focussing on the eccentric mergers has the advantage that we know they must originate from a dynamical or triple channel and that these channels make distinct predictions for the spin-orbit properties. Second, we did not consider the possibility that multiple channels contribute to the formation of eccentric mergers. We expect this to influence our analysis only if they contribute at about the same order of magnitude, in which case in step (ii) of Sec. II a mixture model must be constructed that contains a fraction $0<\beta<1$ of mutually aligned mergers and a fraction $1-\beta$ of uncorrelated ones. Third, we approximated the posterior distribution of $\chi_{\rm eff}$ with a truncated normal distribution, but note that in reality asymmetries may arise (e.g., Ng et al., 2018). However, we do not expect this to significantly change our qualitative results. Furthermore, establishing robust correlations between measurable eccentricity and binary spin remains an open problem in the field. Our analysis is strictly valid in the high–signal-to-noise ratio (SNR) regime; should selection effects due to eccentricity-spin measurements prove to be significant, the methodology presented in this paper can be straightforwardly applied to a higher-purity subset of the catalogue, obtained by imposing a more stringent SNR threshold. While this would increase the required number of detections, it would not alter the qualitative conclusions of the approach.